OpenStudy (anonymous):

If y varies directly as x, find the constant of variation if y = 5 when x = -0.1.

OpenStudy (anonymous):

y = K*x When x = 0.1,-> y = 5 5 = K*(0.1) -> K = 5/(0.1) = 50.